OpenStudy (anonymous):

Charlie was carrying around the steps to solve Square root of 6x plus 4 = 8. On his way through the Math Factory, Charlie dropped them and they mixed with other steps. Put the correct steps in order to solve Square root of 6x plus 4 = 8.

OpenStudy (anonymous):

sqrt(6x+4) 8 (sqrt(6x+4))² = 8² 6x+4 = 64 6x +4 = 64-4 6x = 60 6x/6 = 60/6 x=10